# Exploring Conformational Landscapes Along Anharmonic Low-Frequency Vibrations

### v1

@@ II.2 FRESEAN Mode Analysis of Trajectory Fragments
To analyze equilibrium fluctuations of the peptide, we first analyzed the time evolution of the peptide backbone dihedrals in the unbiased simulation. We then extracted two trajectory fragments of 700 ps (C) and 400 ps (C) length that describe fluctuations within two distinct conformational states of the peptide but no transitions between them. We then performed FRESEAN mode analysis as described in our previous work[46] independently for both trajectory fragments. For the FRESEAN mode analysis[46], we rotated coordinates and velocities of the peptide trajectory into a common frame of reference and computed a matrix of mass–weighted time cross-correlations between velocities for all degrees of freedom and of the peptide (including auto-correlations for ) with a maximum lag time of 2.0 ps:
@@
Here, we focused our analysis on vibrational modes contributing to the VDoS at zero frequency. At zero frequency, the FRESEAN modes (eigenvectors) with the six largest eigenvalues describe rigid-body translations and rotations of the alanine dipeptide (see Figure S1 and Figure S2 of the Supplementary Information, SI). The remaining eigenvectors with non-zero eigenvalues describe vibrations with significant contributions to the zero-frequency spectrum, which use as an indication of anharmonic properties. For both trajectroy fragments (states C and C), we select non-translational and non-rotational FRESEAN modes at zero frequency with the two largest eigenvalues (modes 7 and 8) as CVs for enhanced sampling simulations described below.
@@ II.3 Metadynamics Simulations
For each of the conformational states C and C, we performed an independent enhanced sampling simulation using CVs and starting configurations obtained from the corresponding trajectory fragments of 700 and 400 ps length, respectively. As described above, we selected eigenvectors seven and eight of the cross correlation matrix at zero frequency as CVs in each case, which were combined with the well-tempered (WT) metadynamics simulation protocol[57] as implemented in the PLUMED 2.6.6[72] plugin for GROMACS. Additionally we also performed a third WT-metadynamics simulation using the two dihedral angles as CVs. All metadynamics simulations were run for 50 ns in the NPT ensemble using the same general simulation parameters as described above for the equilibrium simulations unless otherwise noted.
@@ III.1 Vibrational Analysis of Unbiased Trajectories
In Figure 1, we illustrate the two dihedral angles phi () and psi () of alanine peptide (panel A) and plotted their time evolution during a 10 ns unbiased simulation (panel B). It is apparent that rotations around are fast and transitions between favored states at approximately 170 (conformation C) and 0 (conformation C) are frequently observed on the timescale of the simulation. In contrast, rotations around are associated with sufficiently high kinetic barriers that impede sampling in unbiased simulations. Earlier studies show that spontaneous rotations occur only on the s timescale despite the low complexity of the system.[74]
@@
Our goal is to enable the generation of suitable CVs with as little information as possible. We thus limit our vibrational analysis of alanine dipeptide to short fragments of the 10 ns trajectory in Fig. 1B that exclusively describe local fluctuations within a single conformational state, i.e., either C or C. We highlighted two selected trajectory fragments in Fig. 1B that are 700 ps (C) and 400 ps (C) in length.
We then performed FRESEAN mode analysis independently for these two trajectory fragments. First, we plotted the VDoS of the alanine dipeptide for C or C at far-infrared frequencies in Figure 2A, which we computed from velocity time auto correlation functions as defined in eq. 3. While the two conformers exhibit slightly distinct vibrations for wavenumbers greater than 200 cm, they share a broad vibrational band with a peak at 50cm.
@@
Modes 7 and 8 describe low-frequency vibrations with peak frequencies between 50 and 100 cmand non-zero intensities at zero frequency as shown in Figure 5. In panels A and C, we plot the isolated 1D-VDoS for fluctuations along modes 7 and 8, respectively, as observed in both analyzed trajectory fragments for conformational states C and C. The 1D-VDoS are calculated from time auto correlations of projections of weighted velocities () of the peptide atoms on a given FRESEAN mode with components .
As for the FRESEAN mode analysis, weighted velocities are defined in a reference frame that follows the rotations of the peptide during the simulation. We note that apart from the broad main peak, the 1D-VDoS for all four vibrational modes (modes 7 and 8 for C and C) are free from contributions of high-frequency vibrations of the peptide at frequencies outside the range shown in Figure 5. As discussed in our previous work and in contrast to harmonic or quasi-harmonic mode analysis, FRESEAN mode analysis can be used to successfully isolate low-frequency vibrations of molecular systems.[46]
The broad line shape of the peaks in the 1D-VDoS and the intensity at zero frequency (proportional to the eigenvalue of ) can be interpreted as indicators of an anharmonic vibration. We visualized the displacement vectors associated with each vibrational mode in panels B and D of Figure 5, which allow for a qualitative characterization of the nature of the vibrations. For mode 7, in both the C and C conformations, the dominant feature of the displacement vectors are opposing motions of the carbonyl oxygen and amide hydrogen in the N-terminal peptide bond, which implies a rotation of the angle (see Figure 1A). Likewise, displacement vectors for mode 8 in both conformations describe opposing motions of the carbonyl oxygen and amide hydrogen in the C-terminal peptide bond as expected for a rotation of the angle. While each vibrational mode describes a collective motion with contributions from all peptide atoms, both dihedral angles of the peptide clearly have major contributions. The latter is significant because this information did not enter the analysis and we explicitly analyzed trajectory fragments that did not exhibit large amplitude fluctuations of the dihedrals. Hence, the FRESEAN mode analysis identified anharmonic vibrational modes associated with both dihedral angles without prior information purely from local fluctuations. Notably, this was repeated independently for local fluctuations in conformation C and conformation C.
@@ III.2 Enhanced Sampling Simulations along Anharmonic Vibrations
We then tested whether the anharmonic vibrations identified with the FRESEAN mode analysis are suitable CVs in enhanced sampling simulations. Specifically, we tested whether CVs based on these vibrations enable efficient sampling of the conformational space of alanine dipeptide. For this purpose, we performed a set of three independent well-tempered WT-metadynamics simulations as described in the Methods section. Two simulations started from conformations C and C, respectively, and used the corresponding zero frequency FRESEAN modes 7 and 8 as CVs. The third simulation was performed as a reference and used the dihedral angles and as CVs.
Free energy surfaces obtained from the first two simulations, which use the FRESEAN modes of conformations C and C to define the biasing potential, are displayed in Figure 6. Differences between both free energy surfaces are expected, because distinct sets vibrational modes obtained from distinct trajectory fragments corresponding to conformations C and C, as well as distinct reference structures, are used to define the projection of atomic coordinates into the FRESEAN mode space. In Figure 7A, we plotted the free energy surface obtained from the third simulation as a function of the dihedral angles and used as CVs to define the biasing potential in that case. To allow for direct comparisons between all three simulations, we further plotted free energy surfaces in dihedral angle space in panels B and C of Figure 7, which we obtained after unbiasing the metadynamics simulations obtained with FRESEAN modes as CVs. For this purpose, we computed the thermodynamic weight of each trajectory time frame () based on its free energy in vibrational mode space (defined by a projection of displacements relative to a reference structure on FRESEAN modes 7 and 8, denoted in the following as and ).[75]
@@
A notable characteristic of the free energy surfaces in Figure 6 is the small magnitude of free energy barriers between distinct minima in the space described by the anharmonic vibrations of conformations C and C. Several free energy minima are observed that are connected to each other with barriers lower than 10 kJ/mol. A single barrier of approximately 25 kJ/mol is observed in Figure 6A (based on conformations C and its vibrations) for direct transitions between two states indicated as 2 and 3 (more details on this assignment in the following). However, no equivalent barrier is seen for indirect transitions passing through state 1 or in the free energy surfaces shown in Figure 6B (based on conformations C and its vibrations).
The latter is especially striking in comparison to the free energy surface obtained in our reference simulation with the backbone dihedral angles and as CVs (Figure 7A). The latter reproduces multiple previous results in the literature[54, 58, 59] and describes free energy barriers below 10 kJ/mol for transitions between states indicated as 1 and 2 (equivalent to conformations C and C observed in the unbiased simulation trajectory in Figure 1) via rotations of . However, the lowest free energy pathway that connects states 1 and 2 with state 3 in dihedral angle space involves a rotation of with a free energy barrier of 25 kJ/mol.
@@
To identify how free energy minima in dihedral angle space are mapped into the FRESEAN mode space, we determined the average structure of configurations corresponding to states 1, 2 and 3 in Figure 7A. We then computed the root mean-square displacements (RMSDs) relative to these three structures for all structures observed in the metadynamics simulations in FRESEAN mode space based on C and C, respectively. We averaged the RMSDs for all time frames with identical projections in the corresponding FRESEAN mode space and the results are shown in Figure 8. The result allows a direct mapping of the free energy minima observed in dihedral angle space in Figure 7A onto distinct regions of the free energy surfaces observed in the FRESEAN mode space in Figure 6A and B. The latter was used to label the conformational states 1, 2 and 3 from Figure 7A in panels A and B of Figure 6.

### v2

@@ II.2 FRESEAN Mode Analysis of Trajectory Fragments
To analyze equilibrium fluctuations of the peptide, we first analyzed the time evolution of the peptide backbone dihedrals in the unbiased simulation. We then extracted two trajectory fragments of 700 ps (C1) and 400 ps (C2) length that describe fluctuations within two distinct conformational states of the peptide but no transitions between them. We then performed FRESEAN mode analysis as described in our previous work[46] independently for both trajectory fragments. For the FRESEAN mode analysis[46], we rotated coordinates and velocities of the peptide trajectory into a common frame of reference and computed a matrix of mass–weighted time cross-correlations between velocities for all degrees of freedom and of the peptide (including auto-correlations for ) with a maximum lag time of 2.0 ps:
@@
Here, we focused our analysis on vibrational modes contributing to the VDoS at zero frequency. At zero frequency, the FRESEAN modes (eigenvectors) with the six largest eigenvalues describe rigid-body translations and rotations of the alanine dipeptide (see Figure S1 and Figure S2 of the Supplementary Information, SI). The remaining eigenvectors with non-zero eigenvalues describe vibrations with significant contributions to the zero-frequency spectrum, which use as an indication of anharmonic properties. For both trajectroy fragments (states C1 and C2), we select non-translational and non-rotational FRESEAN modes at zero frequency with the two largest eigenvalues (modes 7 and 8) as CVs for enhanced sampling simulations described below.
@@ II.3 Metadynamics Simulations
For each of the conformational states C1 and C2, we performed an independent enhanced sampling simulation using CVs and starting configurations obtained from the corresponding trajectory fragments of 700 and 400 ps length, respectively. As described above, we selected eigenvectors seven and eight of the cross correlation matrix at zero frequency as CVs in each case, which were combined with the well-tempered (WT) metadynamics simulation protocol[57] as implemented in the PLUMED 2.6.6[72] plugin for GROMACS. Additionally we also performed a third WT-metadynamics simulation using the two dihedral angles as CVs. All metadynamics simulations were run for 50 ns in the NPT ensemble using the same general simulation parameters as described above for the equilibrium simulations unless otherwise noted.
@@ III.1 Vibrational Analysis of Unbiased Trajectories
In Figure 1, we illustrate the two dihedral angles phi () and psi () of alanine peptide (panel A) and plotted their time evolution during a 10 ns unbiased simulation (panel B). It is apparent that rotations around are fast and transitions between favored states at approximately 170 ∘ (conformation C1) and 0 ∘ (conformation C2) are frequently observed on the timescale of the simulation. In contrast, rotations around are associated with sufficiently high kinetic barriers that impede sampling in unbiased simulations. Earlier studies show that spontaneous rotations occur only on the s timescale despite the low complexity of the system.[74]
@@
Our goal is to enable the generation of suitable CVs with as little information as possible. We thus limit our vibrational analysis of alanine dipeptide to short fragments of the 10 ns trajectory in Fig. 1B that exclusively describe local fluctuations within a single conformational state, i.e., either C1 or C2. We highlighted two selected trajectory fragments in Fig. 1B that are 700 ps (C1) and 400 ps (C2) in length.
We then performed FRESEAN mode analysis independently for these two trajectory fragments. First, we plotted the VDoS of the alanine dipeptide for C1 or C2 at far-infrared frequencies in Figure 2A, which we computed from velocity time auto correlation functions as defined in eq. 3. While the two conformers exhibit slightly distinct vibrations for wavenumbers greater than 200 cm-1, they share a broad vibrational band with a peak at 50cm-1.
@@
Modes 7 and 8 describe low-frequency vibrations with peak frequencies between 50 and 100 cm-1and non-zero intensities at zero frequency as shown in Figure 5. In panels A and C, we plot the isolated 1D-VDoS for fluctuations along modes 7 and 8, respectively, as observed in both analyzed trajectory fragments for conformational states C1 and C2. The 1D-VDoS are calculated from time auto correlations of projections of weighted velocities () of the peptide atoms on a given FRESEAN mode with components .
As for the FRESEAN mode analysis, weighted velocities are defined in a reference frame that follows the rotations of the peptide during the simulation. We note that apart from the broad main peak, the 1D-VDoS for all four vibrational modes (modes 7 and 8 for C1 and C2) are free from contributions of high-frequency vibrations of the peptide at frequencies outside the range shown in Figure 5. As discussed in our previous work and in contrast to harmonic or quasi-harmonic mode analysis, FRESEAN mode analysis can be used to successfully isolate low-frequency vibrations of molecular systems.[46]
The broad line shape of the peaks in the 1D-VDoS and the intensity at zero frequency (proportional to the eigenvalue of ) can be interpreted as indicators of an anharmonic vibration. We visualized the displacement vectors associated with each vibrational mode in panels B and D of Figure 5, which allow for a qualitative characterization of the nature of the vibrations. For mode 7, in both the C1 and C2 conformations, the dominant feature of the displacement vectors are opposing motions of the carbonyl oxygen and amide hydrogen in the N-terminal peptide bond, which implies a rotation of the angle (see Figure 1A). Likewise, displacement vectors for mode 8 in both conformations describe opposing motions of the carbonyl oxygen and amide hydrogen in the C-terminal peptide bond as expected for a rotation of the angle. While each vibrational mode describes a collective motion with contributions from all peptide atoms, both dihedral angles of the peptide clearly have major contributions. The latter is significant because this information did not enter the analysis and we explicitly analyzed trajectory fragments that did not exhibit large amplitude fluctuations of the dihedrals. Hence, the FRESEAN mode analysis identified anharmonic vibrational modes associated with both dihedral angles without prior information purely from local fluctuations. Notably, this was repeated independently for local fluctuations in conformation C1 and conformation C2.
@@ III.2 Enhanced Sampling Simulations along Anharmonic Vibrations
We then tested whether the anharmonic vibrations identified with the FRESEAN mode analysis are suitable CVs in enhanced sampling simulations. Specifically, we tested whether CVs based on these vibrations enable efficient sampling of the conformational space of alanine dipeptide. For this purpose, we performed a set of three independent well-tempered WT-metadynamics simulations as described in the Methods section. Two simulations started from conformations C1 and C2, respectively, and used the corresponding zero frequency FRESEAN modes 7 and 8 as CVs. The third simulation was performed as a reference and used the dihedral angles and as CVs.
Free energy surfaces obtained from the first two simulations, which use the FRESEAN modes of conformations C1 and C2 to define the biasing potential, are displayed in Figure 6. Differences between both free energy surfaces are expected, because distinct sets vibrational modes obtained from distinct trajectory fragments corresponding to conformations C1 and C2, as well as distinct reference structures, are used to define the projection of atomic coordinates into the FRESEAN mode space. In Figure 7A, we plotted the free energy surface obtained from the third simulation as a function of the dihedral angles and used as CVs to define the biasing potential in that case. To allow for direct comparisons between all three simulations, we further plotted free energy surfaces in dihedral angle space in panels B and C of Figure 7, which we obtained after unbiasing the metadynamics simulations obtained with FRESEAN modes as CVs. For this purpose, we computed the thermodynamic weight of each trajectory time frame () based on its free energy in vibrational mode space (defined by a projection of displacements relative to a reference structure on FRESEAN modes 7 and 8, denoted in the following as and ).[75]
@@
A notable characteristic of the free energy surfaces in Figure 6 is the small magnitude of free energy barriers between distinct minima in the space described by the anharmonic vibrations of conformations C1 and C2. Several free energy minima are observed that are connected to each other with barriers lower than 10 kJ/mol. A single barrier of approximately 25 kJ/mol is observed in Figure 6A (based on conformations C1 and its vibrations) for direct transitions between two states indicated as 2 and 3 (more details on this assignment in the following). However, no equivalent barrier is seen for indirect transitions passing through state 1 or in the free energy surfaces shown in Figure 6B (based on conformations C2 and its vibrations).
The latter is especially striking in comparison to the free energy surface obtained in our reference simulation with the backbone dihedral angles and as CVs (Figure 7A). The latter reproduces multiple previous results in the literature[54, 58, 59] and describes free energy barriers below 10 kJ/mol for transitions between states indicated as 1 and 2 (equivalent to conformations C1 and C2 observed in the unbiased simulation trajectory in Figure 1) via rotations of . However, the lowest free energy pathway that connects states 1 and 2 with state 3 in dihedral angle space involves a rotation of with a free energy barrier of 25 kJ/mol.
@@
To identify how free energy minima in dihedral angle space are mapped into the FRESEAN mode space, we determined the average structure of configurations corresponding to states 1, 2 and 3 in Figure 7A. We then computed the root mean-square displacements (RMSDs) relative to these three structures for all structures observed in the metadynamics simulations in FRESEAN mode space based on C1 and C2, respectively. We averaged the RMSDs for all time frames with identical projections in the corresponding FRESEAN mode space and the results are shown in Figure 8. The result allows a direct mapping of the free energy minima observed in dihedral angle space in Figure 7A onto distinct regions of the free energy surfaces observed in the FRESEAN mode space in Figure 6A and B. The latter was used to label the conformational states 1, 2 and 3 from Figure 7A in panels A and B of Figure 6.